OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar Two resistances \(\sf R_1\) and \(\sf R_2\) are made of different materials. the temperature coefficient of the material of \(\sf R_1~is~\alpha\) and of \(\sf R_2~is~- \beta\). The resistance of the series combination of \(\sf R_1~and~R_2\) will not change with temperature, then ratio of resistance of two wire at 0°C will be ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

@nincompoop

OpenStudy (abhisar):

@abb0t @kropot72

OpenStudy (abhisar):

@Somy

OpenStudy (somy):

and where is the circuit

OpenStudy (abhisar):

There is no circuit

OpenStudy (abhisar):

@iambatman can u help ?

OpenStudy (anonymous):

I would but I'm about to go buy myself a new carpet xD.

OpenStudy (abhisar):

aaaahhh come on it's ur left hand's play XD

OpenStudy (abhisar):

answer is \(\sf \frac{\beta}{\alpha}\) btw

OpenStudy (abhisar):

We got the answers here, but not able to understand http://openstudy.com/updates/53e2690be4b0045ba5512d19

OpenStudy (anonymous):

Is it the coefficient you don't understand?

OpenStudy (abhisar):

yes...

OpenStudy (abhisar):

This line If things don't change with temperature, then \(\alpha - \beta = 0 \Rightarrow \alpha = \beta\)?

OpenStudy (abhisar):

I am not even able to understand how he solved the equation ?

OpenStudy (anonymous):

You have to ignore any changes due to thermal expansion, the resistance of each section changes with temperature, so the total resistance of the materials doesn't change with temperature so that's how he got it basically. (If that makes sense)

OpenStudy (anonymous):

Anyways PK did a good job :P, I'll get back to you later if you're still having trouble.

OpenStudy (abhisar):

Just tell me one thing...if resistance does not change then what is its effect on coefficient ?

OpenStudy (abhisar):

Just tell me how he solved the equation ?

OpenStudy (abhisar):

This one \[R_1 + R_2 = r_1 + r_2 + \Delta T \left(r_1 \alpha - r_2 \beta\right)\]\[\Rightarrow r_1 \alpha = r_2\beta\]And blah blah. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

There's a dependance of the resistance for the series combination of resistance at To, do you know the formula?

OpenStudy (abhisar):

no ! What is it ?

OpenStudy (anonymous):

I think you should give this a read before you attack these sort of problems haha :P http://hyperphysics.phy-astr.gsu.edu/hbase/electric/restmp.html

OpenStudy (anonymous):

It's pretty neat stuff

OpenStudy (abhisar):

aaahhh..i know this...:D just not able to how he deduced the last line \(\sf r_1 \alpha =r_2\beta\)

OpenStudy (abhisar):

*able to understand

OpenStudy (anonymous):

It simplifies to that because delta T is the same for each material

OpenStudy (anonymous):

Idk if I'm confusing you more or if it's helping XD

OpenStudy (abhisar):

Hmmm...i think i'll have to give some more time on it.

OpenStudy (anonymous):

Maybe I'll give you an example which is some what similar but a bit simpler and it will help you understand.

OpenStudy (abhisar):

yeah..that will be great !

OpenStudy (anonymous):

Alright, well here is a problem: Two wires have the same cross-sectional area and are joined end to end to form a single wire. One is tungsten, which has a temperature coefficient of resistivity of \[\alpha \] 0.0045 (C degrees)^-1. The other is carbon, for which \[\alpha \]= -0.0005 (C degrees)^-1. The total resistance of the composite wire is the sum of the resistances of the pieces. The total resistance of the composite does not change with temperature. What is the ratio of the lengths of the tungsten and carbon sections? Ignore any changes in length due to thermal expansion. Yeah I suck at LaTeX, but I'm too lazy to make it fancy.

OpenStudy (abhisar):

It's ok...i can understand it without latex too :)

OpenStudy (abhisar):

How will u solve it ?

OpenStudy (anonymous):

Well ok it's like the same process, so the total resistance of the composite wire doesn't change... \[(R_t)_0+(R_c)_0 = R_t+R_c\] so left side is at room temp, and right side at temp T

OpenStudy (anonymous):

The temperature dependence of the resistance of the wire of resistance \[R_0\] at temperature \[T_0\] is given by \[R=R_0=[1+\alpha(T-T_0)]\] \[(R_t)_0+(R_c)_0=(R_t)_0(1+ \alpha_t \Delta T)+(R_c)_0(1+\alpha_c \Delta T)\] and for this delta T is same for each wire, amg the similarities with the other question XD

OpenStudy (anonymous):

This is cutting into my carpet buying time but I will finish!

OpenStudy (abhisar):

so \(\sf \alpha _t=\alpha _c\) ?

OpenStudy (anonymous):

\[(R_t)_0 \alpha _t=-(R_c) \alpha _c\] oh btw the subscript t = tungsten and subscript c = carbon So with this expression you can find the ratio of the resistances. And once that's known, you can easily find the ratio of the lengths of the sections using the equation \[L = \frac{ RA }{ p }\] \[\frac{ (R_t)_0 }{ (R_c)_0 }=-\frac{ \alpha_c }{ \alpha_t }=-\frac{ -0.0005 C degrees ^{-1} }{ 0.0045 C degrees ^{-1} }\]

OpenStudy (anonymous):

\[= \frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

\[\frac{ L_t }{ L_c }=\frac{ (R_0A/p)_t }{ (R_0 A/p)_c }=\frac{ (R_t)_0 }{ (R_c)_0 }\ \left( \frac{ p_c }{ p_t } \right)\] ok i lied this was harder than your question lmao.

OpenStudy (abhisar):

;(

OpenStudy (anonymous):

Did that help at all ? :S

OpenStudy (abhisar):

Yes it did !

OpenStudy (anonymous):

Alright, cool haha, I think it just requires practice and eventually you'll understand, just do a few problems related to these kinds of things.

OpenStudy (abhisar):

Yes..i tdo think so !

OpenStudy (abhisar):

Thanx for ur tremendous Help !

OpenStudy (anonymous):

Alright AB, I gtg now, take care and good luck!! ^_^

OpenStudy (anonymous):

No problemo

OpenStudy (abhisar):

Now i think u can go buy ur Carpet xD

OpenStudy (anonymous):

Yeah haha xD

OpenStudy (abhisar):

Byeee ! :) tcu2

OpenStudy (abhisar):

ahaaan ! I think i really got it.......... :D Thank you @iambatman ...wish i could gift uh a carpet XD

OpenStudy (abhisar):